OpenStudy (anonymous):

write the slope-intercept form of the equation of the line passing through (4, -6) and parallel to the line y= 8x-2

OpenStudy (aripotta):

what do we know about parallel lines?

OpenStudy (anonymous):

I've kinda forgotten everything to be honest.

OpenStudy (aripotta):

ok so parallel lines can't intersect each other. so they both have the same slope. since we know that the line goes through point (4,-6) and the slope must be 8, we can use point-slope form: y - y1 = m(x - x1)

OpenStudy (aripotta):

so plug in your info and then convert it into slope-intercept form

OpenStudy (anonymous):

You found a pro alex xD

OpenStudy (anonymous):

what would i use for y and y1 and x and x1

OpenStudy (aripotta):

y and x stay the same. y1 is the y part of the point and x1 is the x part of the point. so y1 = -6 and x1 = 4

OpenStudy (anonymous):

and m would be 8 right?

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (aripotta):

wait you're not done yet tho cuz you have to convert it into slope-intercept form before you submit it

OpenStudy (anonymous):

This is such simple math but I love watching you break it down for alex @AriPotta

OpenStudy (anonymous):

#Proatwork

OpenStudy (anonymous):

so since y1 would be -6 and the equation is y-y1 would it be y-(-6) and then the rest of the equation?

OpenStudy (aripotta):

y - (-6) = 8(x - 4) y + 6 = 8(x - 4)

OpenStudy (aripotta):

when you subtract a negative, it turns into addition

OpenStudy (aripotta):

but you're not done yet since your answer has to be in slope-intercept form

OpenStudy (anonymous):

I'm not sure if I've done the equation right so far. i started with y-(-6)= 8(x-4) then i switched it to y+6= 8x-32 and subtracted 6 from each side and got y=8x - 38

OpenStudy (aripotta):

that's exactly right :)

OpenStudy (anonymous):

is that all I have to do?

OpenStudy (aripotta):

yep.

OpenStudy (anonymous):

okay thank you! :)

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)